What are the missing numbers for this problem? 120,60,180,90, _,_,405

Answer by AnlytcPhil(1806)



120,60,180,90, _,_,405
120
120÷2 = 60
60×3 = 180
180÷2 = 90
90×3 = 270
270÷2 = 135
135×3 = 405